Quantum Mechanics At what angles do the diffracti Help plz
Answer #1


a) torque = I x alpha

torque is r x F .

(1.5 * 100 * sin90) - ( 0.8 *150 *sin40) = (25 *2^2 / 3) * alpha

alpha = 2.19 rad/s^2


b) angular momentum = I w

         = (25 * 2^2 / 3) * 25 = 833.33 kg m^2 /s


c) KE = Iw^2 /2 = ( 25* 2^2 / 3) * (25^2) / 2 = 10416.67 J